40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10289     Entry time: Tue Jul 29 19:00:40 2014
Author: Harry 
Type: Update 
Category: General 
Subject: Weekly Plan (7.29.14) 

 The Past Week

In the past week, I have improved the coupling in the fiber testing setup on the SP table to up to ~45%

I also measured the input/output modes of the fiber with collimators.

Manasa, Q and I have designed, and redesigned a setup to measure Polarization Extinction Ratio introduced by fibers.

I have also partially assembled the box that will hold the frequency counters and RPi for FOL.

Today (Tuesday) I measured waists of PSL and AUX, at dumped light from the SHG's for use in designing coupling telescopes for FOL.

Next Week

In the next week, I will design and couple light from PSL and AUX (Y arm) into fibers for use in testing FOL.

Once that's done, I will continue testing fiber characteristics, starting with Polarization Extinction Ratio.

Items Needed

Power cord for Raspberry Pi (ordered)

AD9.5F collimator adapter (ordered)

 

ELOG V3.1.3-